Show that fo = 1/2lT/u is homogenous
1
Expert's answer
2020-11-10T07:09:06-0500

We have an equation:


"f_0=\frac{1}{2l}\sqrt{\frac T\mu}."

This expression describes the frequency of a string with length "l", mass per unit length "\mu", under tension "T." This expression only applies to homogenous string, i.e., strings of constant "\mu". If mass per unit length is not constant, the string cannot form nodes and antinodes and it is impossible to compute its frequency.